OpenStudy (anonymous):

A 0.400 ice puck, moving east with a speed of 5.52 , has a head-on collision with a 0.950puck initially at rest. Assuming a perfectly elastic collision, what will be the speed of the 0.400 object after the collision? What will be the speed of the 0.950 object after the collision? Please provide a detailed solution, Im am confused

OpenStudy (anonymous):

so in physics, everytime the word, "collision" is in the problem, use 'conservation of momentum' formula. does that help, or should I help more? ^_^

OpenStudy (anonymous):

I know i need to get it to the quadratic but I am struggling how to get there

OpenStudy (anonymous):

mm... i'm not too sure of needing to get it into a quadratic, but I can show you how to set it up mv + mv = mv + mv so the left side of the = is the 'before collision' and the right sid of the = is the 'after collision' so, we now have 0.4(5.52) + 0.95(0) = mv + mv 0.4(5.52) + 0.95(0) = 0.4(v) + 0.95(v) making sense now?

OpenStudy (anonymous):

i'm logging off, i'll give you one more step: 0.4(5.52) + 0.95(0) = 0.4(v) + 0.95(v) becomes 0.4(5.52) + 0.95(0) = v(0.4 + 0.95) solve for v ^_^

OpenStudy (anonymous):

oh thats where I went wrong thanks for your help